How many pounds of peanuts costing $2.00 a pound should be mixed with a 4 pounds of cashews costing $4.50 a pound to obtain a mixture costing $3.00 a pound?
Answer by checkley71(8403)

4*450+x*200=(4+x)300
1800+200x=1200+300x
-100x=-600
x--600/-100
x=6 pounds of $4.50 cashews. thus we need 4+6=10 pou8nds of peanuts.
proof
4*450+6*200=10*300
1800+1200=3000
3000=3000